consider the following frequency table representing the scores on a test
class. Frequency
40-44. 3
45-49. 4
50-54. 12
55-59. 4
60-64. 15

determine the lower class boundary for the third class
Answer #1

Solution :

=> The distance from the upper limit of the first class to the lower limit of the second class is 45 - 44 = 1

=> Half this distance is = 0.5

=> The lower boundary of third class is calculated by subtracting half of the gap value from the third class lower limit

=> Third class lower boundary = 50 - 0.5 = 49.5